OpenStudy (anonymous):

Two cyclists, one averaging 12 miles per hour and the other 28 miles per hour, start from the same town at the same time. If they travel in opposite directions, after how long will they be 88 miles apart?

OpenStudy (anonymous):

Can you please tell me how you got that?

OpenStudy (cwrw238):

combined speed = 12 + 28 = 40 mph time to travel 88 miles = 88/40 = 2.2 hours 2 hours 12 minutes

OpenStudy (anonymous):

ok, i ll try both ways :)

OpenStudy (cwrw238):

1.1 hours cannot be right because in that time cyclyst travelling at 12 mph covers just over 12 miles and other covers about 30 miles so thats about 42 miles apart

OpenStudy (cwrw238):

as they are travelling in opposite directions the speed they are travelling apart = 12+28 = 40 mph

OpenStudy (anonymous):

I think cwrw238 is correct! (yay)

OpenStudy (cwrw238):

yep - thats right so the time taken is 88/40 hours

OpenStudy (anonymous):

Either way, thanks for helping me! Muchly appreciated :)

OpenStudy (anonymous):

sorry sorry @cwrw238 is correct..